40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Cryo Lab eLog  Not logged in ELOG logo
Message ID: 2005     Entry time: Thu Apr 5 15:28:38 2018
Author: Zach 
Type: Update 
Category: SiFi - ringdown 
Subject: Newest hole cantilevers RT Q = ~80k 

I performed ringdowns on the most recent batch of hole cantilevers (see CRYO:2004). This is the first set of ringdowns done with the newer clamp design (i.e., the ones that go into the main experiment) and with the precise clamp alignment procedure.

 

Both cantilevers had a frequency of 94 Hz. I made two ringdown measurments on each one, with a full removal and reclamp between trials.

Sample "2/27/2018 #1A1" exhibited a Q of 83k on the first trial, and again 83k on the second trial. This is near the thermoelastic limit for this specimen, and the reproducibility is very nice.

Sample "2/27/2018 #1A2" exhibited a Q of 88k on the first trial, but only 74k on the second trial.

Therefore, it seems that while the new clamps are capable of being very robust, it is still possible for clamping effects to come into play.

 


Also, I found the odd triangular-connector breakout cable that I need to interface with the feedthrough on the smaller cryostat! After tearing the Cryo Lab up looking for it, I found it in the EE Shop :-)

ELOG V3.1.3-